OpenStudy (anonymous):

Two ropes, AD and BD, are tied to a peg on the ground at point D. The other ends of the ropes are tied to points A and B on a flagpole, as shown below: Two ropes, AD and BD, are tied to a peg on the ground at point D. The other ends of the ropes are tied to points A and B on a flagpole. Angle ADC measures 60 degrees and angle BDC measures 30 degrees. The length of DC is 5 multiplied by square root of 3. Angle ADC measures 60° and angle BDC measures 30°. What is the distance between the points A and B on the flagpole?

OpenStudy (anonymous):

can anyone help me out with this

OpenStudy (anonymous):

the answer choices are 10, 15, 20, and 25

OpenStudy (campbell_st):

is C the base of the flagpole..?

OpenStudy (campbell_st):



OpenStudy (anonymous):

i posed a picture of the triangle

OpenStudy (campbell_st):

so find the length AC using the tan ratio and an angle of 60 then find BC, again using the tan ratio and an angle of 30 the the distance AB is AC - BC hope it helps

OpenStudy (anonymous):

thank you! wait is the answer 10ft?

OpenStudy (campbell_st):

I don't know I haven't done the calculation

OpenStudy (anonymous):

can you walk me through it

OpenStudy (campbell_st):

ok... so do you know the exact value of tan(30) and tan(60)..?

OpenStudy (anonymous):

tan60 is 1.73205080757

OpenStudy (anonymous):

tan30 is 0.57735026919

OpenStudy (campbell_st):

here is the start\[\tan(60) = \frac{AC}{5 \sqrt{3}}~~~then~~~AC = 5 \sqrt{3} \times \tan(60)\] and \[\tan(30) = \frac{BC}{5\sqrt{3}}~~~~~then~~~~BC = 5\sqrt{3} \times \tan(30)\] hope it helps

OpenStudy (campbell_st):

last step AB = AC - BC

OpenStudy (anonymous):

i got 20ft?

OpenStudy (campbell_st):

I think you need to check the calculations

OpenStudy (anonymous):

ok let me check

OpenStudy (anonymous):

ohhhhh ok so now i see what i did wrong is the answer 15?

OpenStudy (campbell_st):

well AC is 15 now calculate BC

OpenStudy (anonymous):

so bc is 5 right?

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

that would mean the answer is 10 right?

OpenStudy (campbell_st):

correct

OpenStudy (anonymous):

THANK SO MUCH !!!!!!!